# Critic-Guided Reinforcement Unlearning in Text-to-Image Diffusion

### v2

@@ 5.3 Concept Removal
We used a trained CLIP-based classifier as our reward function, utilizing the ”openai/clip-vit-base-patch32” model. The classifier outputs probabilities across the 20 object classes. We used complement probability of the target class, scaled into 0-10 range, as a reward signal.
With the reward function established, we proceeded to train our per-timestep critic network using the same CLIP model for consistency.
@@
To further validate our approach, we conducted a detailed comparison between CGRU and DDPO on the ”Cats” class. We use this comparison to, once again, experimentally verify the validity of our theoretical claims about the convergence benefits of CGRU compared to standard RL fine-tuning, in the field of machine unlearning. Both methods were trained using identical experimental conditions, with the only difference being CGRU’s incorporation of the per-timestep critic. Figure 2 presents the comparative results, demonstrating CGRU’s superior performance in concept suppression.
Figure 2: Concept suppression comparison between CGRU and DDPO on the ”Cats” class. CGRU demonstrates superior performance in reducing unwanted object generation.
@@ 6 Evaluation
We present comprehensive evaluation results comparing CGRU against state-of-the-art unlearning methods on the UnlearnCanvas benchmark. Our evaluation encompasses both quantitative metrics and qualitative assessments to demonstrate the effectiveness of our approach.
@@ 6.1 Quantitative Results
We evaluate CGRU’s performance using established metrics from the UnlearnCanvas benchmark Zhang et al. (2024b), computed using the ViT classifier provided by the creators of UnlearnCanvas. Our evaluation focuses on three key metrics that capture different aspects of unlearning effectiveness:
Unlearning Accuracy (UA) measures the proportion of samples generated from prompts containing the target concept that are not correctly classified by the concept detector. In-domain Retain Accuracy (IRA) quantifies the model’s ability to correctly generate and classify samples containing other concepts from the same domain, ensuring that unlearning does not severely harm unrelated capabilities. We also measure image quality through Fréchet Inception Distance (FID) Heusel et al. (2018) to ensure that unlearning does not compromise the visual fidelity of generated images.
@@
Method UA ($\uparrow$) IRA ($\uparrow$) FID ($\downarrow$)
ESD Gandikota et al. (2023) $92.15\%$ $55.78\%$ $65.55$
FMN Zhang et al. (2024b) $45.64\%$ $90.63\%$ $131.37$
UCE Gandikota et al. (2024) $\underline{94.31}\%$ $39.35\%$ $182.01$
CA Kumari et al. (2023) $46.67\%$ $90.11\%$ 54.21
SalUn Fan et al. (2024) $86.91\%$ $\mathbf{96.35}\%$ $61.05$
SEOT Li et al. (2024) $23.25\%$ $\underline{95.57}\%$ $62.38$
SPM Lyu et al. (2024) $71.25\%$ $90.79\%$ $\underline{59.79}$
EDiff Wu et al. (2024) $86.67\%$ $94.03\%$ $81.42$
SHS Wu and Harandi (2024) $80.73\%$ $81.15\%$ $119.34$
SAeUron cywiński2025saeuroninterpretableconceptunlearning $78.82\%$ $95.47\%$ $62.15$
CGRU $\mathbf{95.55}\%$ $78.47\%$ $98.43$
Table 1 presents the comparative results on object unlearning tasks. The baseline results are taken from cywiński2025saeuroninterpretableconceptunlearning, and our method is added for comparison. We provide a Pareto frontier visualization of the UA-IRA trade-off in Appendix E. CGRU demonstrates strong performance in unlearning accuracy (95.55%), achieving the best result among all methods. While the retain accuracy (78.47%) is competitive, it shows that CGRU effectively suppresses target concepts while maintaining reasonable capability to generate other objects.
The results validate our theoretical framework, showing that the per-timestep critic approach enables more effective unlearning compared to existing methods that rely on parameter editing or global penalties. CGRU’s balanced performance across all metrics demonstrates its ability to achieve the core objective of machine unlearning: removing specific concepts while preserving overall model capabilities.
@@ 6.2 Qualitative Results
To visualize the progression of concept forgetting during training, we compare generated images from different checkpoints between CGRU and DDPO for a representative object class. Figure 3 shows the evolution of model behavior as training progresses, demonstrating the differences in how each method handles concept suppression.
Figure 3: Progression of concept forgetting during training for CGRU and DDPO. Each row shows generated images from different training checkpoints (default, early, mid, late), demonstrating how each method handles concept suppression while preserving overall image quality.
The qualitative comparison reveals key differences between the approaches: (i) Both methods start with clear presence of target objects in the default (untrained) state, (ii) DDPO shows aggressive concept suppression but suffers from significant degradation in image quality and coherence, resulting in poor retain performance, (iii) CGRU demonstrates more balanced forgetting, effectively suppressing target concepts while maintaining natural-looking images and preserving the ability to generate other objects. Additional qualitative results are provided in Appendix E.
7 Discussion & Limitations
An interesting observation is that methods achieving high unlearning accuracy tend to exhibit lower retain accuracy (see Figure 5), as seen across different approaches in the literature. This trade-off appears to be inherent to diffusion unlearning approaches. We believe investigating the fundamental causes of this tension is a substantial research direction that warrants a dedicated study, which we leave for future work.
CGRU outperforms DDPO, validating that per-timestep critics provide more effective optimization signals than sparse, end-of-trajectory rewards. The timestep-aware critic architecture enables precise credit assignment throughout the denoising process. The FiLM-enhanced critic proves particularly effective at learning meaningful value estimates across different stages of the denoising process.
While our experiments demonstrate effectiveness on object removal tasks, the generalizability of CGRU to other types of unlearning remains to be investigated.
@@ 7 Discussion & Limitations
Our method requires task-specific reward function and critic design and the performance depends on their quality. Training requires substantial computational resources, making it computationally expensive for rapid prototyping or large-scale deployment.
Unlike parameter editing methods that provide theoretical guarantees about concept removal, our reinforcement learning approach offers limited formal guarantees about unlearning completeness. The success relies on learned components subject to approximation errors.
@@ 8 Conclusion
We presented Critic-Guided Reinforcement Unlearning (CGRU), a framework that treats denoising as a sequential decision process with per-timestep feedback. Our timestep-aware critic provides dense signals throughout the denoising trajectory, enabling effective credit assignment and reduced variance compared to sparse rewards. CGRU achieves 95.55% unlearning accuracy on the UnlearnCanvas benchmark, demonstrating superior performance over existing methods.
Future work should explore adaptive reward functions, investigate robustness against adversarial prompts, and extend the framework to other unlearning types. Our code and evaluation scripts are available at Anonymous (2025) to facilitate reproducibility and encourage further research in RL-based diffusion unlearning.
@@ Appendix A Proof of Advantage-Weighted Policy Gradient for Diffusion
Let $x_{0}$ denote the final sample generated from a conditional diffusion model given context $c$. Sampling proceeds by drawing $x_{T}\sim\mathcal{N}(0,I)$ and iteratively sampling $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t},c)$ to obtain trajectory $\tau=(x_{T},x_{T-1},\ldots,x_{0})$ with terminal reward $r(x_{0},c)$. Viewing denoising as a multi-step MDP with states $s_{t}=(x_{t},c,t)$ and actions $a_{t}=x_{t-1}$, the policy is $\pi_{\theta}(a_{t}|s_{t})\equiv p_{\theta}(x_{t-1}|x_{t},c)$ and rewards are zero at all $t>0$ and $r(x_{0},c)$ at $t=0$.
@@ A.2 Objective and the DDPO score-function gradient
The RL objective for diffusion is to maximize expected terminal reward under the distribution of samples induced by the reverse process:
$J(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0}|c)}[r(x_{0},c)]$ (15)
@@ A.4 Baseline-augmented (advantage) gradient
Plugging $f(\tau)=r(x_{0},c)$ and then averaging over $c\sim p(c)$ gives Eq. equation 16.
The next step is introduction of the critic. We will integrate it into the loss, prove that the change does not bias the original objective and the impact of such changes. Fix an arbitrary baseline function $b(x_{t})$ that depends only on the state $x_{t}$ (it may also depend on $c$ and $t$, but it must not depend on the sampled action $x_{t-1}$). Start from the gradient expression and add zero in the form “BBB-B” where $B$ is the baseline-weighted term:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot r(x_{0},c)\right]+\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot b(x_{t})\right]$
$-\underbrace{\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot b(x_{t})\right]}_{\;=:\;B\;-\;B\;}$
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot(r(x_{0},c)-b(x_{t}))\right]+B,$ (19)
@@
$B:=\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot b(x_{t})\right]$
@@
$B=\int p(c)\int\left[\sum_{t=1}^{T}b(x_{t})\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\right]p_{\theta}(\tau|c)d\tau dc.$
@@
$\int b(x_{t})\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)p_{\theta}(\tau|c)d\tau=\int\left[b(x_{t})\left(\int p_{\theta}(x_{t-1}|x_{t},c)\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)dx_{t-1}\right)\right]p_{\theta}(\text{rest}|x_{t},c)d(\text{rest})dx_{t}.$

### v3

@@ 5.3 Concept Removal
We trained a CLIP-based classifier to serve as our reward function, utilizing the ”openai/clip-vit-base-patch32” model with classification head. The classifier outputs probabilities for the 20 object classes. We used the complement probability of the target class, scaled in the 0-10 range, as a reward signal.
With the reward function established, we warm-start the per-timestep critic network using the same CLIP model for consistency and continue updating it online during diffusion model fine-tuning.
@@
5.4 NSFW Content Removal.
To test whether CGRU extends beyond object-level erasure to safety-critical concept removal, we apply it to suppress nudity/NSFW generations and evaluate on a standardized prompt suite, specifically - I2P benchmark Schramowski et al. (2023). This setting is challenging because the target concept is highly abstract and can be elicited by diverse prompts.
We use a public binary NSFW detector Falconsai (2023) (a fine-tuned vision transformer) as the reward function. Concretely, given an image $x$, the detector outputs a probability of NSFW content $p_{\text{nsfw}}(x)$; we set the reward as $r(x)=1-p_{\text{nsfw}}(x)$, so higher reward corresponds to safer generations. We warm-start a per-timestep critic with the same backbone architecture and FiLM conditioning as in our CGRU setup for concept removal and keep updating it online to predict the expected terminal reward from intermediate noisy latents.
To obtain a diverse distribution of explicit and borderline content, we draw training images from the NSFW Detection Dataset by DeepGHS deepghs (2023), which contains five categories (drawing, hentai, neutral, porn, sexy). We train the critic on a random 10% split and perform policy optimization on a disjoint 20% split. Since RL fine-tuning requires text prompts, we convert images into prompts by re-captioning each image using LLaVA-1.5-7B Liu et al. (2023). We use a safety-aware instruction (Appendix G.1) that explicitly requests a literal description and asks the model to include NSFW or suggestive elements when present; this reduces prompt ambiguity and makes the unlearning objective better aligned with the safety concept being removed. We further assess robustness to adversarial prompt attacks on the UnlearnDiff nudity benchmark and its accompanying UnlearnDiffAtk procedure (adversarial prompt-rewriting attack) Zhang et al. (2024d). Table 1 reports the Attack Success Rate (ASR) before and after applying UnlearnDiffAtk; CGRU achieves 0.00 pre-ASR and 9.67 post-ASR.
Table 1: Attack Success Rate (ASR) before and after UnlearnDiffAtk Zhang et al. (2024d) on UnlearnDiff nudity prompts (lower is better).
Method Pre-ASR Post-ASR
FMN 88.03 97.89
SPM 54.93 91.55
SAFREE 26.06 85.59
UCE 21.83 79.58
ESD 20.42 76.05
RECE 13.38 75.42
MACE 9.10 74.57
AdvUn 7.75 21.13
SalUn 1.41 11.27
SHS 0.00 7.04
Erasediff 0.00 2.11
CGRU (ours) 0.00 9.67
5.5 Ablation: Impact of Timestep Awareness
To validate the importance of our timestep-aware design, we compare the FiLM-conditioned critic against a standard CLIP-based critic without timestep conditioning. Both models are evaluated on their ability to predict the final reward from intermediate noisy latents. Since the reward derives from a 20-class concept classifier, we frame this as a classification task over reward levels; test samples are drawn uniformly across timesteps to cover the full noise spectrum.
The timestep-aware critic achieves 43.25% accuracy and 58.52% macro precision, compared to 29.75% and 34.47% for the baseline (see Figure 10 in Appendix E). This substantial improvement confirms that FiLM-based timestep conditioning is essential for accurate value estimation from noisy states, particularly in the earlier stages of diffusion where the signal-to-noise ratio is low.
@@ 6 Evaluation
We evaluate CGRU on (i) object concept removal using the UnlearnCanvas benchmark Zhang et al. (2024b) and (ii) safety-critical nudity suppression using the Inappropriate-to-Proper (I2P) benchmark from Safe Latent Diffusion Schramowski et al. (2023). We report both quantitative metrics and qualitative evidence to characterize the erasure–utility trade-off.
@@ 6.1 Quantitative Results
We first evaluate CGRU on UnlearnCanvas Zhang et al. (2024b) using the ViT classifier provided by the benchmark. We focus on three metrics that capture different aspects of unlearning: Unlearning Accuracy (UA) measures the proportion of samples generated from prompts containing the target concept that are not correctly classified by the concept detector. In-domain Retain Accuracy (IRA) quantifies the model’s ability to correctly generate and classify samples containing other concepts from the same domain, ensuring that unlearning does not severely harm unrelated capabilities. We also measure image quality through Fréchet Inception Distance (FID) Heusel et al. (2018) to ensure that unlearning does not compromise the visual fidelity of generated images.
@@
Method UA ($\uparrow$) IRA ($\uparrow$) Average ($\uparrow$) FID ($\downarrow$)
ESD Gandikota et al. (2023) $92.15\%$ $55.78\%$ $73.97\%$ $65.55$
FMN Zhang et al. (2024b) $45.64\%$ $90.63\%$ $68.14\%$ $131.37$
UCE Gandikota et al. (2024) $\underline{94.31}\%$ $39.35\%$ $66.83\%$ $182.01$
CA Kumari et al. (2023) $46.67\%$ $90.11\%$ $68.39\%$ 54.21
SalUn Fan et al. (2024) $86.91\%$ $\mathbf{96.35}\%$ $\mathbf{91.63}\%$ $61.05$
SEOT Li et al. (2024) $23.25\%$ $\underline{95.57}\%$ $59.41\%$ $62.38$
SPM Lyu et al. (2024) $71.25\%$ $90.79\%$ $81.02\%$ $\underline{59.79}$
EDiff Wu et al. (2024) $86.67\%$ $94.03\%$ $\underline{90.35\%}$ $81.42$
SHS Wu and Harandi (2024) $80.73\%$ $81.15\%$ $80.94\%$ $119.34$
SAeUron cywiński2025saeuroninterpretableconceptunlearning $78.82\%$ $95.47\%$ $87.15\%$ $62.15$
CGRU $\mathbf{95.37}\%$ $1.64\% 88$$8.50\%$ $68.43$
Table 2 summarizes object unlearning results. Baseline numbers are taken from the recent consolidated benchmark in cywiński2025saeuroninterpretableconceptunlearning; CGRU is evaluated with the same UnlearnCanvas protocol. CGRU attains the highest UA (95.37%), while trading off IRA.
Table 3: Nudity unlearning evaluation on the I2P benchmark. CLIPScore is calculated for the NSFW experiment model only. The best result for each metric is highlighted in bold.
Method Armpits Belly Buttocks Feet Breasts (F) Genitalia (F) Breasts (M) Genitalia (M) Total CLIPScore ($\uparrow$)
FMN Zhang et al. (2024b) 43 117 12 59 155 17 19 2 424 30.39
CA Kumari et al. (2023) 153 180 45 66 298 22 67 7 838 31.37
AdvUn Zhang et al. (2024c) 8 0 0 13 1 1 0 0 28 28.14
Receler Huang et al. (2024) 48 32 3 35 20 0 17 5 160 30.49
MACE Lu et al. (2024) 17 19 2 39 16 0 9 7 111 29.41
CPE Lee et al. (2025b) 10 8 2 8 6 1 3 2 40 31.19
UCE Gandikota et al. (2024) 29 62 7 29 35 5 11 4 182 30.85
SLD-M Schramowski et al. (2023) 47 72 3 21 39 1 26 3 212 30.90
ESD Gandikota et al. (2023) 32 30 2 19 27 3 8 2 123 30.21
SAeUron cywiński2025saeuroninterpretableconceptunlearning 7 1 3 2 4 0 0 1 18 30.89
CGRU 2 6 7 6 17 0 0 2 40 30.27
SD v1.4 148 170 29 63 266 18 42 7 743 31.34
SD v2.1 105 159 17 60 177 9 57 2 586 31.53
Figure 2: UA–IRA trade-off (Pareto frontier). Each point is a method from Table 2. CGRU achieves the strongest erasure (UA) and lies above the baseline trend for high-erasure regimes, highlighting it as a Pareto-competitive choice when prioritizing concept removal.
Across methods, there is a clear tension between suppressing the target concept (high UA) and preserving in-domain capability for related concepts (high IRA), consistent with prior observations in diffusion unlearning. Table 2 shows that CGRU achieves the strongest UA (95.37%) while maintaining a solid IRA (81.64%), yielding the third-best overall Average score (88.50%). While retention-optimized baselines such as SalUn and SAeUron attain higher IRA, they do so with weaker UA; conversely, other high-UA erasure methods (e.g., ESD/UCE) sacrifice substantially more IRA. Figure 2 highlights this regime: CGRU sits above the baseline UA–IRA trend among common erasure methods (ESD/UCE/SHS/SPM/SEOT), making it the best choice when UA must be prioritized. Second, we evaluate nudity suppression on I2P using NudeNet vladmandic (2023) with a 0.6 confidence threshold (following Schramowski et al. (2023)); detections are aggregated per category and summarized as a total. On I2P, CGRU reduces the total nudity detections to 40 (Table 3), substantially improving over concept erasure baselines such as ESD and UCE, and approaching the strongest reported methods such as SAeUron. Importantly, we also report CLIPScore to monitor prompt adherence: while aggressive safety unlearning can reduce alignment to the input text, CGRU maintains acceptable semantic fidelity. We note that CGRU’s CLIPScore (30.27) is lower than retention-focused baselines; this reflects a known trade-off where aggressive nudity suppression can reduce prompt adherence, particularly when prompts themselves contain suggestive language. Balancing erasure with explicit alignment preservation (e.g., via multi-objective rewards) is a promising direction for future work. Overall, this result suggests that the critic-guided RL framework extends beyond object-level erasure to a more abstract, safety-critical concept (nudity), using an evaluation protocol that is independent of the training reward.
@@ 6.2 Qualitative Results
Qualitatively, we visualize how concept evidence decays as RL fine-tuning progresses. Figure 3 shows generations from intermediate checkpoints for a representative prompt under CGRU. Beyond simply removing a target concept, we observe that the model often “fills in” the scene in a visually plausible way: the removed object may be replaced by a different semantically compatible subject (e.g., another object or a human figure), blended into the surrounding environment, or expressed only through non-diagnostic background elements. In other words, forgetting typically manifests as a natural re-composition of the image rather than a degenerate collapse into noise or low-quality artifacts.
Figure 3: Progression of concept forgetting during training for CGRU. Images are sampled from different training checkpoints (default, early, mid, late), illustrating the gradual suppression of the target concept.
To avoid overloading the main paper with large image grids, we defer additional qualitative progressions (Figures 67), per-class UA/IRA (Table 4), and per-class misclassification matrices (Figure 4) to Appendix B. We also report a gradient-variance analysis supporting the stability claim (Appendix B.4) and details on the FID/CLIPScore calculation in the Appendix F.
7 Limitations & Future Work
While our experiments demonstrate effectiveness on both object concept removal (UnlearnCanvas) and safety-critical nudity suppression (I2P), the generalizability of CGRU beyond these settings remains to be investigated.
@@ 7 Limitations & Future Work
CGRU relies on task-specific automated reward signals (e.g., concept detectors or NSFW classifiers) and a critic that must learn to predict them from noisy latents; performance is therefore bounded by the fidelity and coverage of these rewards. Moreover, RL fine-tuning is computationally expensive, which limits rapid prototyping and large-scale deployment. Unlike some parameter editing methods that offer stronger formal guarantees, our reinforcement learning approach provides limited guarantees about unlearning completeness. The success relies on learned components subject to approximation errors. Additionally, the effectiveness of CGRU depends on the quality of the reward model; a poorly specified detector may lead to reward hacking (e.g., generating images that fool the detector without true concept removal). We mitigate this by following established benchmarks and using publicly available detector architectures, but ensemble rewards or human evaluation could provide additional robustness. As with any output-level optimization, CGRU may re-route generation pathways rather than fully erase latent concept representations. The qualitative evidence (Figure 3) suggests the model redistributes visual content in a semantically plausible way rather than collapsing to noise, which is desirable for image quality but leaves open the question of whether deeper concept removal occurs. More generally, unlearning may hold under the evaluated prompt distribution while still allowing regrowth under distribution shift or adaptive prompting; while we include an adversarial prompt evaluation for nudity (Table 1), robustness to stronger attacks remains open. Finally, we do not include an explicit retention loss or retain-set training; retention is an emergent property of localized parameter updates and the base model prior, and can degrade under distribution shift.
Future work.
Promising directions include improving robustness under prompt variation and adaptive attacks, and developing more principled multi-objective formulations that explicitly trade off erasure and retention. It will also be valuable to extend CGRU beyond object removal and nudity suppression (e.g., to styles, identities, or other safety domains). Evaluating on newer diffusion backbones (e.g., SDXL, SD3) is another important step; we focus on Stable Diffusion 1.5 primarily due to the computational cost of RL training and because it remains the common evaluation standard for state-of-the-art baselines. Finally, we believe it is promising to study where and how unlearning manifests inside the model using interpretability and parameter/representation analysis methods (e.g., weight/activation attribution, probing, sparse autoencoders, or related mechanistic analyses), with the goal of making unlearning more targeted and easier to diagnose.
@@ 8 Conclusion
We presented Critic-Guided Reinforcement Unlearning (CGRU), an RL framework for diffusion unlearning that treats reverse denoising as a sequential decision process and augments policy-gradient updates with a timestep-aware critic. By conditioning value estimation on intermediate noisy latents, CGRU provides dense learning signals across the trajectory, improving credit assignment and stabilizing training compared to sparse terminal-reward optimization. Empirically, CGRU achieves strong concept suppression on UnlearnCanvas, reaching 95.37% unlearning accuracy while maintaining reasonable retention and image quality, and extends to safety-critical nudity suppression on I2P. We complement aggregate benchmark numbers with per-class diagnostics and misclassification analyses in the appendix, and include supporting ablations and stability evidence. Our code and evaluation scripts are available at Anonymous (2026) to facilitate reproducibility and follow-up research on RL-based diffusion unlearning.
@@ Appendix A Proof of Advantage-Weighted Policy Gradient for Diffusion
Let $x_{0}$ denote the final sample generated from a conditional diffusion model given context $c$. Sampling proceeds by drawing $x_{T}\sim\mathcal{N}(0,I)$ and iteratively sampling $x_{t-1}\sim p_{\theta}(x_{t-1}|x_{t},c)$ to obtain trajectory $\tau=(x_{T},x_{T-1},\ldots,x_{0})$ with terminal reward $r(x_{0},c)$. Viewing denoising as a multi-step MDP with states $s_{t}=(x_{t},c,t)$ and actions $a_{t}=x_{t-1}$, the policy is $\pi_{\theta}(a_{t}|s_{t})\equiv p_{\theta}(x_{t-1}|x_{t},c)$ and rewards are zero at all $t>0$ and $r(x_{0},c)$ at $t=0$. The MDP can be written as
$s_{t}$ $=(x_{t},c,t)$ (1)
$a_{t}$ $=x_{t-1}$ (2)
$\pi_{\theta}(a_{t}|s_{t})$ $=p_{\theta}(x_{t-1}|x_{t},c)$ (3)
$P(s_{t+1}|s_{t},a_{t})$ $=\delta_{x_{t-1}}\otimes\delta_{t-1}$ (4)
$R(s_{t})$ $=\begin{cases}r(x_{0},c)&\text{if }t=0\\ 0&\text{otherwise.}\end{cases}$ (5)
Here the context $c$ is fixed for the entire episode and therefore omitted from the transition kernel notation. The initial state distribution is $\rho_{0}(s_{0})=p(x_{T})\otimes\delta_{T}$ with $x_{T}\sim\mathcal{N}(0,I)$. The objective is the expected terminal reward under the induced sample distribution:
$J(\theta)=\mathbb{E}_{c\sim p(c)}\mathbb{E}_{x_{0}\sim p_{\theta}(x_{0}|c)}[r(x_{0},c)].$ (6)
@@ A.4 Baseline-augmented (advantage) gradient
Plugging $f(\tau)=r(x_{0},c)$ and then averaging over $c\sim p(c)$ gives Eq. (7).
The next step is introduction of the critic. We will integrate it into the loss, prove that the change does not bias the original objective and the impact of such changes. Fix an arbitrary baseline function $b(x_{t},c,t)$ that depends only on the state $x_{t}$ (it may also depend on $c$ and $t$, but it must not depend on the sampled action $x_{t-1}$). Start from the gradient expression and add zero in the form “BBB-B” where $B$ is the baseline-weighted term:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot r(x_{0},c)\right]+\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot b(x_{t},c,t)\right]$
$-\underbrace{\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot b(x_{t},c,t)\right]}_{\;=:\;B\;-\;B\;}$
@@
$\nabla_{\theta}J(\theta)=\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot(r(x_{0},c)-b(x_{t},c,t))\right]+B,$ (10)
@@
$B:=\mathbb{E}_{c,\tau}\left[\sum_{t=1}^{T}\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\cdot b(x_{t},c,t)\right]$
@@
$B=\int p(c)\int\left[\sum_{t=1}^{T}b(x_{t},c,t)\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)\right]p_{\theta}(\tau|c)d\tau dc.$
@@
$\int b(x_{t},c,t)\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)p_{\theta}(\tau|c)d\tau=\int\left[b(x_{t},c,t)\left(\int p_{\theta}(x_{t-1}|x_{t},c)\nabla_{\theta}\log p_{\theta}(x_{t-1}|x_{t},c)dx_{t-1}\right)\right]$
$p_{\theta}(\text{rest}|x_{t},c)d(\text{rest})dx_{t}.$